### 02 Jan 2024 v1

# Has your pretrained model improved? a multi-head posterior based approach

Abstract
The emergence of pretrained models has significantly impacted from Natural Language Processing (NLP) and Computer Vision to relational datasets. Traditionally, these models are assessed through fine-tuned downstream tasks. However, this raises the question of how to evaluate these models more efficiently and more effectively. In this study, we explore a novel approach where we leverage the meta features associated with each entity as a source of worldly knowledge and employ entity representations from the models. We propose using the consistency between these representations and the meta features as a metric for evaluating pretrained models. Our method’s effectiveness is demonstrated across various domains, including models with relational datasets, large language models and images models.
1 Introduction
Pre-training on large models is becoming increasingly common in various machine learning applications, thanks to the growing amount of user-generated content. This is evident in areas such as Natural Language Processing (NLP) with models like GPT (Generative Pretrained Transformer), and in the vision-language domain with models like CLIP. Typically, the effectiveness of these models is evaluated using downstream tasks. However, these can be relatively costly if all tasks need to be performed. Is there a more efficient, yet simple method to evaluate these models?
Entity representations as known as embeddings, generated from these models, can be utilized directly or indirectly by downstream tasks and can also be fine-tuned as needed. The associated meta features with these embeddings can be considered as the model’s foundational knowledge of the world it’s learning from. This could be the class category for image data or semantic and syntactic information for words. Despite having the same meta features, embeddings differ across models. The degree of consistency between the embeddings and meta features can therefore serve as a performance metric for model evaluation.
The embedding space is complex and challenging to interpret or explain. Despite extensive efforts to decipher it, the intricacies of the embedding space go beyond mere linear interpretability as some research suggests. In this research, we hypothesize that the embeddings reside within a manifold space where Euclidean distance is not an appropriate metric for gauging the similarity between two embeddings. Meta features are capable of grouping these embeddings into clusters, we assume each forming a sub-manifold space. When the clusters are fine granulated enough, we can assumingly use a Gaussian distribution to approximate each cluster. Taken together, these clusters constitute a mixture of Gaussian distributions. By calculating the posterior probabilities of the entities, we can calculate the consistency of the meta features and embeddings.
In this study, we introduce a unique approach to evaluate the performance of pretrained models. Specifically, we:
  1. 1.
    Adopt a distinct perspective towards the model. Instead of focusing solely on downstream performance, we place greater emphasis on the quality of the entities’ representations that the model can generate.
  2. 2.
    Consider the features associated with the entity representations as the benchmark to assess their quality. We hypothesize that the meta features can partition the embedding space into distinct clusters. The quality of these clusters can be evaluated using the posterior probability of Gaussian mixture models.
  3. 3.
    While there are multiple methods to interpret this meta feature spaces, we present a tree-based approach as an example that uses meta features to segment entities into clusters.
  4. 4.
    Test the effectiveness of our proposed method on a variety of datasets across different domains, ranging from recommendation-based models to language models and image models. We present both qualitative and quantitative evidence to demonstrate the approach’s efficacy.
2 Algorithm Framework
For a given domain, we first collect a large size of entities with rich meta features. Then for any given pretrained model, we can generate an embedding dataset denoted as $\mathbf{X}={\mathbf{x}_{1},...,\mathbf{x}_{N}}$, where each $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and $1\leq i\leq N$. Here, $N$ represents the number of entities and $d$ signifies the dimension of the embeddings. Simultaneously, we can produce a corresponding feature set $\mathbf{F}={\mathbf{f}_{1},...,\mathbf{f}_{N}}$. Each feature $\mathbf{f}_{i}$ comprises both categorical and numerical features. We convert numerical features into categorical ones for consistency. The primary objective is to examine the consistency between these two datasets, $\mathbf{X}$ and $\mathbf{F}$.
In the simplest scenario where $\mathbf{f}_{i}$ has only one feature, a straightforward method of segmentation is to form clusters based solely on these features, with each category creating its own cluster. However, when $\mathbf{f}_{i}$ has $m$ features, and each feature has $k_{j}$ categories, the number of combinations becomes $\prod_{j=1}^{m}k_{j}$. This results in a significantly larger number of clusters. We will postpone the discussion on finding the best split to a later session. In the upcoming discussion, we will assume that we already have a split criterion for the dataset $\mathbf{X}$.
2.1 Posterior Based Embedding Evaluating Metric
We aim to evaluate the effectiveness of different models that generate these embeddings to determine the best one. The splitting criteria $\mathbf{S}$ can divide the entities into a group of clusters $C_{1},C_{2},...,C_{n}$, with each entity belonging to a single cluster, where $n$ is number of clusters. To evaluate the quality of the cluster, we adopt a posterior based method.
In the context of GMM, it is assumed that the data is generated from a combination of multiple Gaussian distributions. Each component of this mixture corresponds to one of the distinct clusters within the dataset. For any given data point within a particular cluster, the formula for calculating its posterior probability in the GMM framework can be expressed as follows:
$P(\theta=k|\mathbf{x}_{k})=\frac{P(\mathbf{x}_{k}|\theta=k)P(\theta=k)}{\sum_{j=1}^{m}P(\mathbf{x}_{k}|\theta=j)P(\theta=j)},$ (1)
where $\mathbf{x}_{k}$ represents every point in cluster $k$. $P(\theta=k)$ = number of pionts in cluster $k$ / $N$.
To assess the quality of embeddings $\mathbf{X}$ within the context of a splitting $\mathbf{S}$, we compute the overall evaluation metric $P_{\mathbf{S}}^{\mathbf{X}}$ by averaging the log probabilities of all the embeddings across all clusters. This metric provides an assessment of the quality of the embeddings. We call it the average of log posterior.
$ALP_{\mathbf{S}}^{\mathbf{X}}=\frac{1}{N}\sum_{k=1}^{m}\sum_{\mathbf{x}_{i}\in C_{k}}\log P(\theta=k|\mathbf{x}_{i})$ (2)
One challenge with the aforementioned formula is its high sensitivity to outliers. A single outlier could lead to an extremely large value for $P_{\mathbf{S}}^{\mathbf{X}}$. To mitigate the impact of such outlier entities, we implement a clipping mechanism for embeddings with very small posterior probabilities. Specifically, if $P(\theta=k|\mathbf{x}_{k})$ is less than $k/N*\varepsilon$, we exclude the entity from the $P_{\mathbf{S}}^{\mathbf{X}}$ computation.
Another challenge arises when the embeddings exist in large dimensions. If the number of embeddings in each cluster is smaller than the embedding dimension, this leads to a rank-deficient covariance. To address this, we propose a multi-head solution. In this approach, we randomly select $v$ dimensions and evaluate based on these dimensions. This process is repeated multiple times and the average results are used. This concept is inspired by the random forest algorithm (Breiman, 2001). Additionally, we apply the routine regularization approach, i.e., adding $\epsilon$$I$ to the covariance matrix. The value of $\epsilon$ is decided in the following manner.
$\epsilon=max(\lambda_{k}/(10D)\lambda_{0},1e{-}8),$ (3)
where $D$ is the dimensionality of the embeddings and $\lambda_{i}$ are the eigenvalues of the covariance matrix (sorted decreasingly by their magnitude). $k$ is the minimum value that satisfies $\frac{\Sigma_{i=0}^{k}\lambda_{i}}{\Sigma_{i=0}^{D}\lambda_{i}}>99.99\%$.
2.2 One Meta Feature Based Clustering
In the simplest scenario, where the feature vector $\mathbf{f}_{i}$ consists of only one feature, a clear and intuitive approach to segmentation is to form clusters based solely on these features. This method is straightforward as it capitalizes on the inherent characteristics of the data. Each unique category within the data forms its own distinct cluster, effectively grouping similar items together. The consistency of these clusters with the embeddings can serve as a measure of the quality of the embeddings. However, it’s important to note that extending this approach to accommodate more than two meta features is not as straightforward.
2.3 Meta features + representation based segmentation
Inspired by EmbeddingTree algorithm, we can construct the a tree based on the entities and all the leaf nodes are the the finally clusters, specifically: we first convert non-binary categorical features into binary ones by asking yes-no questions regarding to each of their categorical values and get the binary feature sets: $\mathbf{G}=\{\mathbf{g}_{1},...,\mathbf{g}_{N}\}$ ($\mathbf{g}_{i}\in\{0,1\}^{q}$, $1\leq i\leq N$), $q$ denote the total number of converted binary features.
With the processed data, we describe the EmbeddingTree algorithm with details in Algorithm 1. We iterate through the $q$ features (line 6) and evaluate them based on the splitting criteria described in Section 2.4 to pick out the best feature for splitting (line 8-10), using the feature’s binary value (line 11-13). The above procedure is executed recursively (line 15-16) until the splitting criterion ($\Theta$), e.g., the number of entities per tree node or the tree depth, is no longer satisfied (line 2). With the given embedding and feature data, the whole procedure is deterministic.
Algorithm 1 Build an EmbeddingTree
1:procedure BuildTree($[\mathbf{X},\mathbf{F}],q,\Theta$)
2:     if $\Theta\hskip 2.84526ptis\hskip 2.84526ptnot\hskip 2.84526ptsatisfied$  then
3:         return LeafNode([$\mathbf{X},\mathbf{F}$])
4:     else
5:         $max\_t\leftarrow-\infty$
6:         for $k\in\{1,...,q\}$ do
7:              $t=Embedding-MAP([\mathbf{X},\mathbf{F}^{k}])$
8:              if $t>max\_t$ then
9:                  $bestFea=k$
10:                  $max\_t=t$
11:     $[\mathbf{X},\mathbf{F}]_{left}=\{\mathbf{x}\in\mathbf{X}|\mathbf{F}_{bestFea}==0\}$
12:
13:     $[\mathbf{X},\mathbf{F}]_{right}=\{\mathbf{x}\in\mathbf{X}|\mathbf{F}_{bestFea}==1\}$
14:
15:     Children.Left = $BuildTree([\mathbf{X},\mathbf{F}]_{left},q,\Theta)$
16:     Children.Right = $BuildTree([\mathbf{X},\mathbf{F}]_{right},q,\Theta)$
17:     return Children
2.4 2-GMM Splitting with Maximum A Posteriori Estimation (MAP)
One critical component of Algorithm 1 is the criterion used to select the best splitting feature. The criterion is computed based on the approximate MAP for GMMs inspired by (Zheng et al., 2023).
We assume the embedding can be modeled as two mixture Gaussians. The expectation-maximization (EM) algorithm is used to jointly estimate all the parameters and latent variables. The latent variables, $z_{i,j}$, denote the probability that sample $i$ is in cluster $j$. With $N$ as the number of observations and $J$ be the number of Gaussian clusters (in this case, $J=2$), $z=\{z_{1,1},z_{1,2},...,z_{N,J-1},z_{N,J}\}$, the complete likelihood (including the latent variables) is:
$P(\mathbf{x},\mu,\Sigma,w,z)=\prod_{i=1}^{N}\prod_{j=1}^{J}\{w_{j}\mathcal{N}(\mathbf{x}_{n};\mu_{j},\Sigma_{j}^{2})\}^{z_{i,j}},$ (4)
where $\mu$ is the mean vectors and $\Sigma$ is covariance matrix of the Gaussians.
To find the best binary feature that splits the embedding and forms the best GMM, we go through every feature. Each candidate binary feature splits the embeddings into two clusters, each is then formulated as a Gaussian. For each feature, suppose the first $s$ embeddings have feature value $F^{k}=0$ and the rest $N-s$ embeddings have feature value $F^{k}=1$. We estimate the weights, means, and variances for both clusters using the maximum likelihood estimation (MLE).
[…]
In other words, rather than the soft clustering of GMM, our algorithm performs a hard clustering. Thus, if $\mathbf{x}_{i}$ is in cluster $j$, then $z_{i,j}=1$ and $z_{i,j^{\prime}}=0$ for all $j\neq j^{\prime}$. Given this approximation, the likelihood can be obtained by summing over the $z$:
$P(\mathbf{x},\mu,\Sigma,w)=\sum_{z}\prod_{i=1}^{N}\prod_{j=1}^{J}\{w_{j}\mathcal{N}(\mathbf{x}_{n};\mu_{j},\Sigma_{j}^{2})\}^{z_{i,j}}$ (5)
Note that $z_{(i\in(0,s],j=1)}=z_{(i\in[s+1,N),j=2)}=1$ and $z_{i,j}=0$, otherwise, the above equation can be simplified to:
$P(\mathbf{x},\mu,\Sigma,w)=\prod_{i=1}^{s}w_{1}\mathcal{N}(\mathbf{x}_{i};\mu_{1},{\Sigma_{1}}^{2})\prod_{i=s+1}^{N}w_{2}\mathcal{N}(\mathbf{x}_{i};\mu_{2},{\Sigma_{2}}^{2}).$ (6)
We can treat each split feature as another random variable $\theta$, to choose the best split feature, we want to maximum $P(\mathbf{x},\mu,\Sigma,w,\theta)$, in another words, we want to find $\theta$ gives the largest $P(\mathbf{x},\mu,\Sigma,w)$.
2.5 Finding the Best Splitting Point
For simplicity, we only consider $\theta$ as the random variable we want to estimate, by injecting the prior information into the formula, we can treat each splitting feature with different weight. By applying Maximum A Posteriori Estimation (MAP), we can formulate the problem as following:
$P(\theta_{i}|\mathbf{x})=\frac{P(\mathbf{x}|\theta_{i})P(\theta_{i})}{\sum_{j=1}^{q}P(\mathbf{x}|\theta_{j})P(\theta_{j})},$ (7)
where $q$ is number of possible splits.
By plugging (3) into (4), we can get
$P(\theta_{i}|\mathbf{x})=\frac{\prod_{k=1}^{s}w_{1}\mathcal{N}(\mathbf{x}_{k};\mu_{1},{\Sigma_{1}}^{2},\theta_{i})\prod_{k=s+1}^{N}w_{2}\mathcal{N}(\mathbf{x}_{k};\mu_{2},{\Sigma_{2}}^{2},\theta_{i})p(\theta_{i})}{\sum_{j=1}^{q}\prod_{k=1}^{s}w_{1}\mathcal{N}(\mathbf{x}_{k};\mu_{1},{\Sigma_{1}}^{2},\theta_{j})\prod_{k=s+1}^{N}w_{2}\mathcal{N}(\mathbf{x}_{k};\mu_{2},{\Sigma_{2}}^{2},\theta_{j})p(\theta_{j})}.$ (8)
Plugging in the estimates for all the parameters and taking the log of $P(\theta_{i}|\mathbf{x})$,we can get
$\log\hat{P}(\theta_{i}|\mathbf{x})=\sum_{i=1}^{s}[\log\hat{w_{1}}+\log\mathcal{N}(\mathbf{x}_{i};\hat{\mu_{1}},\hat{\Sigma_{1}}^{2})]+\sum_{i=s+1}^{N}[\log\hat{w}_{2}+\log\mathcal{N}(\mathbf{x}_{i};\hat{\mu_{2}},\hat{\Sigma_{2}}^{2})]+\log p(\theta_{i})\\ -\log(\sum_{j=1}^{q}\prod_{k=1}^{s}w_{1}\mathcal{N}(\mathbf{x}_{k};\hat{\mu}_{1},{\hat{\Sigma}_{1}}^{2},\theta_{j})\prod_{k=s+1}^{N}w_{2}\mathcal{N}(\mathbf{x}_{k};\hat{\mu}_{2},{\hat{\Sigma}_{2}}^{2},\theta_{j})p(\theta_{j})).$ (9)
By applying this formula, we can apply the prior knowledge on the importance of the feature to find the split that maximize $\log\hat{P}$.
2.6 Embedding Comparison based on the same splitting criteria
If we have two sets of embeddings, $\mathbf{X_{A}}=\{\mathbf{x_{A}}_{1},\ldots,\mathbf{x_{A}}_{N}\}$, and $\mathbf{X_{B}}=\{\mathbf{x_{B}}_{1},\ldots,\mathbf{x_{B}}_{N}\}$, both trained on the same dataset but using different models, denoted as models $A$ and $B$, where ($\mathbf{x_{A}}_{i},\mathbf{x_{B}}_{i}\in\mathbb{R}^{p}$, $1\leq i\leq N$), we can generate two corresponding splitting criteria, $\mathbf{S_{A}}$ and $\mathbf{S_{B}}$. The objective now is to assess and compare the quality of these two sets of embeddings. Let’s represent $ALP_{\mathbf{S}^{\mathbf{A}}}^{\mathbf{X}^{\mathbf{A}}}$ as $ALP_{\mathbf{A}}^{\mathbf{A}}$ for embeddings $\mathbf{X}_{\mathbf{A}}$ and splitting criteria $\mathbf{S}_{\mathbf{A}}$. Given two sets of embeddings, $\mathbf{X_{\mathbf{A}}}$ and $\mathbf{X_{\mathbf{B}}}$, along with two corresponding splitting criteria, $\mathbf{S_{\mathbf{A}}}$ and $\mathbf{S_{B}}$, we can define four metrics: $ALP_{\mathbf{A}}^{\mathbf{A}}$, $ALP_{\mathbf{B}}^{\mathbf{B}}$, $ALP_{\mathbf{A}}^{\mathbf{B}}$, and $ALP_{\mathbf{B}}^{\mathbf{A}}$. We need to fix the splitting criteria to do clustering, so a proper comparison should be between $ALP_{\mathbf{A}}^{\mathbf{A}}$ and $ALP_{\mathbf{A}}^{\mathbf{B}}$ or between $ALP_{\mathbf{B}}^{\mathbf{A}}$ and $ALP_{\mathbf{B}}^{\mathbf{B}}$.
3 Related Work
3.1 Pretrained models
Large Language Models (LLMs): In recent years, significant strides have been made in the realm of Natural Language Processing (NLP), particularly with the advent of the transformer architecture. Attention-based language models such as BERT (Kenton & Toutanova, 2019), GPT (Brown et al., 2020), XLNet (Yang et al., 2019), and T5 (Raffel et al., 2020) have raised the bar in various language benchmarks. Alongside these developments, a plethora of pre-training and fine-tuning algorithms have been devised to enhance the performance of these transformer models. As these models grew in size, the data-driven nature and scaling characteristics of the transformer architecture became evident. These critical findings paved the way for the creation of large language models (LLMs), including LLaMa 2 (Touvron et al., 2023) with 7-70 billion parameters, BLOOM (Workshop et al., 2022) with 176 billion parameters, and GPT4 (OpenAI, 2023) with an astounding 1.7 trillion parameters. These LLMs demonstrate impressive emergent capabilities, such as solving mathematical equations and analyzing articles, competencies not seen in prior smaller language models. These breakthroughs signify the remarkable progress made in this field.
Vision-Language Pre-training (VLP): With the rapid expansion of model capacity and computational resources, the input to deep neural networks has evolved beyond single modality such as text or image. Vision-Language Pre-training (VLP) was introduced to bridge the gap between different modalities, effectively harnessing cross-modality information from both images and text. Leveraging the successful pre-training and fine-tuning paradigm prevalent in NLP, VLP models have demonstrated exceptional performance in complex vision-language tasks. These tasks include image captioning, visual question answering, and visual reasoning. Among the existing studies, a noteworthy contribution is the CLIP (Radford et al., 2021) model, which employs the concept of contrastive learning to align images and text. CLIP simultaneously trains an image encoder and a text encoder on millions of image-text pairs collected from the internet. The resulting encoders have demonstrated impressive performance on downstream tasks due to their zero-shot classification capability.
Pretrained Dual-Transformers(PDT) for Bipartite Graphs: PDT (Dai et al., 2023)focuses on learning contextual knowledge from a user-content interaction dataset, which is depicted as a bipartite graph. The study identifies two key contexts in the graph: user-side and content-side. The goal of learning from these contexts is framed as two contrastive learning tasks and is applied to a recommendation task. Evaluations on two large popular datasets reveal that PDT outperforms baselines in six metrics.
3.2 Embedding Evaluation Metric
Several significant studies have focused on the evaluation of embeddings in Natural Language Processing (NLP). Schnabel et al. Schnabel et al. (2015) conducted a comprehensive evaluation of word embeddings using both intrinsic and extrinsic measures, focusing on specific tasks and downstream applications. Baroni et al. Baroni et al. (2014) compared count-based and predictive models for word representation, using tasks like semantic and syntactic analogies, word similarity, and concept categorization for evaluation. Conneau et al. Conneau et al. (2018) introduced the Cross-lingual Natural Language Inference (XNLI) corpus as a benchmark for multilingual sentence embeddings evaluation. Niven and Kao Niven & Kao (2019) proposed a new evaluation method for embeddings based on their ability to comprehend logical reasoning in natural language. Zhang et al. Zhang et al. (2020) presented BERTScore, an evaluation metric for text generation models based on BERT embeddings that correlates well with human judgment.
Several works have been proposed to evaluate embeddings between different models or various epochs of the same model (Cui et al., 2022; Ding et al., 2021). One such method is GULP (Boix-Adsera et al., 2022), which ensures consistent control over the difference in prediction performance between two representations for regularized linear prediction tasks. GULP exhibits desirable structural properties, including the triangle inequality and invariance under orthogonal transformations. Recently,  (Gurnee & Tegmark, 2023) demonstrated that Language Models (LLMs) have the capacity to learn linear representations of space and time. These representations are consistent across various entities and exhibit a degree of prompt-independence. The study also suggests that individual neurons can exhibit high sensitivity to these features. Furthermore, the research employed out-of-sample $R^{2}$ as a metric for linear probes, substantiating the hypothesis that larger models tend to perform better.
4 Experimental Analysis
In this experimental session, we initially carry out experiments on a synthetic dataset to verify the effectiveness of our proposed algorithm. Following this, we further evaluate the results in three areas: the MovieLens dataset (Harper & Konstan, 2015) for relational models, spatial datasets (Gurnee & Tegmark, 2023) for large language models, and the Robustness library (Engstrom et al., 2019a; b; Santurkar et al., 2019; 2020) for image models.
4.1 Synthetic Dataset: Gaussian Mixture Model (GMM) of Ten Gaussian Distributions
To validate the efficacy of our propose posterior-based embedding evaluation metric, as outlined in Eq. (2), we designe an experiment that encompass three scenarios, each with $10$ clusters. These clusters are either perfectly separated, partially overlapping, or perfectly overlapping, and are all generated using a Gaussian Mixture Model. Figure 1 presents the results from these scenarios. As anticipated, the Average of Log Posterior (ALP) scores for the 10 perfectly separated Gaussian Distributions was 00, and the accuracy of the clusters assigned from the posterior matrix was $100\%$. In the case of 10 partially overlapping Gaussian Distributions, the ALP score was $-0.3285$, and the accuracy of the clusters assigned from the posterior matrix was $86.96\%$. Similarly, for the 10 perfectly overlapping Gaussian Distributions, the ALP score was $-0.9372$, and the accuracy of cluster assignment was $57.34\%$.
(a) Perfectly separated clusters
(b) Partially overlapping clusters
(c) Perfectly overlapping clusters
Figure 1: Illustration on $2d$ synthetic dataset consisting of 10 gaussian distributions that are perfectly separated, partially overlapping, and perfectly overlapping
4.2 Moive Lens Dataset for Relational
MovieLens-25M consists of 25,000,095 reviews made by 162,541 reviewers on 59,047 movies. We compare with following models to generate the movie embeddings. Word2vec (Mikolov et al., 2013), PDT (Dai et al., 2023) and SASRec (Kang & McAuley, 2018). From the Word2Vec model, we generate two distinct types of embedding representations, specifically w2v_single and w2v_combine. In the case of w2v_single, we create a sequence of movies for each reviewer’s review, sort it based on time, and then employ the Word2vec model to learn the movie embeddings. On the other hand, w2v_combine not only includes the sequences of movies for reviewers but also incorporates sequences of reviewers for each movie and reviewer/movie pairs as part of the training data. For both SASRec and PDT, we generate sequences of nine movies for each reviewer. Additionally, for PDT we also generate sequences of reviewers for each movie, as this is a necessary input for PDT. SASRec is trained using BPR loss, while PDT is trained using two contrastive losses. Both PDT and SASRec are contextualized embeddings, while w2v_single and w2v_combine are static embeddings. We employ two clustering techniques. The first approach involved clustering by single Meta features, such as year and genre. We also apply Embedding tree based method to generate the tree for both year and genre features and use the leaf nodes as clusters.
4.2.1 Movie Lens Dataset: Clustering by Year
In this section, we have evaluated and compared four kinds of embedding representations that were trained on the MovieLens Dataset. These were trained at various iteration levels, we use the “year” feature as labels to cluster the embeddings. As illustrated in Figure 2(a), the PDT and SRSREC embedding performed better than all embeddings across all iterations, as seen in the mean of average of log posteriors plot. During the early stages of training from iteration 1 to 16, w2v_combine outperformed w2v_single. However, in the later stages from iteration 16 to 50, w2v_single superseded w2v_combine.
(a) Mean of average of log posteriors
(b) Mean of accuracy of assigned clusters
Figure 2: Mean of average of the log posterior and accuracy on the MovieLens dataset by clustering on year.
As depicted in the mean of accuracy of assigned clusters plot of Figure 2(b), PDT and SRSREC demonstrates a consistent and stable performance over all other types of embeddings across all iterations. Generally, w2v_single exceeds the performance of w2v_combine. This suggests that contextualized embeddings, specifically PDT and SRSREC, most effectively encode year information and remain stable across all iterations. Also, w2v_single demonstrates superior encoding of year information compared to w2v_combine.
4.2.2 Movie Lens Dataset: Clustering by Genre
By using the genre feature, we create clusters with the genre features as labels. We then compute and report the mean of average of log posteriors and the mean of accuracy of assigned clusters. These findings are presented in Figure 3. Contrary to the consistent pattern observed with year features as labels, the genre features do not exhibit a similar consistency. From Figure 3(a), it’s noticeable that the PDT embedding generally outperforms both SASRec and w2v_single over all iterations. Furthermore, SASRec surpasses w2v_single from the 1st to the 40th iteration, after which they both plateau with similar scores. Between the 12th and 36th iterations, w2v_combine is observed to outperform PDT. Moving to the mean accuracy of assigned clusters plot (Figure 3(b)), it’s evident that PDT consistently outperforms all other embedding types across all iterations. Generally, w2v_combine surpasses both SASRec and w2v_single, except for the first and third iterations where SASRec exceeds w2v_combine.
(a) Mean of average of log posteriors
(b) Mean of accuracy of assigned clusters
Figure 3: Sum of the log posterior on the MovieLens dataset by clustering on genre.
4.2.3 Movie Lens Dataset: Clustering with Tree Leaf Nodes on Genre and Year as the Meta Features
We explore the case where we use the tree leaf nodes from the embedding tree that is constructed with both year and genre as meta features. The mean average of log posteriors of assigned clusters are calculated and reported, as shown in Figure 4(a). The PDT and SASRec embeddings consistently surpass other embeddings throughout all iterations. However, we notice that w2v_combine surpasses w2v_single from the 1st to the 26th iteration, but w2v_single overtakes w2v_combine from the 26th to the 50th iteration. The mean accuracy of assigned clusters, illustrated in Figure 4(b), clearly shows that the PDT and SASRec embedding exhibit a steady and consistent increase in performance compared to all other embeddings across all iterations. This is followed by w2v_single, which generally surpasses w2v_combine.
(a) Mean of average of log Posteriors
(b) Mean of accuracy of assigned clusters
Figure 4: Sum of the log posterior on the movie lens dataset by clustering with tree leaf nodes.
All the above experiment results suggest that the contextualized embeddings SASRec and PDT are more effective at capturing the semantic and structural relationships in the input data compared to their static embedding counterparts, namely w2v_single and w2v_combine, which meet our expectations.
4.3 Experiments with Embedding from Llama-2 Models
In this session, we implement the proposed concept in the field of large language models and use Llama-2 as an example.
Models: The released Llama-2 models (Touvron et al. (2023)) have three versions with different model sizes and embedding dimensionalities. These details have been included in Table 2.
Datasets: We use the Llama-2 models to generate embeddings for three spatial datasets introduced by (Gurnee & Tegmark, 2023). The details of these datasets are shown in Table 2. Each instance of the datasets is a word/phrase for a location (e.g., the name of city, a university, or a place of interest) and based on the spatial location as meta feature, we generate the clusters for each dataset, i.e., based on the continent, state, and borough of the instances from the World Place, USA Place, and NYC Place dataset, respectively.
Table 1: Details of the Llama-2 models.
Model #head #layer #dim
Llama-2-7b 32 32 4096
Llama-2-13b 40 40 5120
Llama-2-70b 64 80 8192
Dataset #clusters #samples cluster
World Place 8 39585 continent
US Place 49 29997 state
NYC Place 7 19838 borough
Table 1: Details of the Llama-2 models.
Table 2: Details of the three spatial datasets.
Embedding Quality: We feed the three datasets into the three Llama-2 models and get their activations/embeddings at each layer of the models. Specifically, we obtain $32$, $40$, and $80$ sets of embeddings for the three sets of models (as the three models have $32$, $40$ and $80$ layers, respectively). For each set of embeddings, we use our proposed method to compute the posterior of individual instance falling into the correct cluster.
In Fig. 5 and 6, In this section, we showcase the quality of the embeddings across three datasets and models. Each plot’s x-axis represents the percentage of layers pertaining to the respective Llama-2 models, while the y-axis signifies the mean of average log posteriors and accuracy. The results reveal a noticeable upward trend in embedding quality from the initial to the final layers. In the USA_Place and NYC_Place datasets, the green lines, representing the larger model (Llama-2-70b), achieve the highest posteriors and accuracy. This suggests that the Llama-2-70b model encodes location information more effectively than the other models. This conclusion holds true for the initial and final interactions of the World_Place dataset, but the consistency wavers during the middle interactions.
Figure 5: Embedding quality over model layers (average posterior). The dimensions are divided into subsets, each comprising 128 dimensions.
Figure 6: Embedding quality over model layers (accuracy). The dimensions are divided into subsets, each comprising 128 dimensions.
Figure 7: Embedding quality over model layers (average posterior). The dimensions are divided into subsets, each comprising 128 dimensions. The covariance matrix for each subset is diagonalized for regularization.
Figure 8: Embedding quality over model layers (accuracy). The dimensions are divided into subsets, each comprising 128 dimensions. The covariance matrix for each subset is diagonalized for regularization.
5 Experiment analysis of CLIP Embeddings on the BREEDS hierarchy
In this section we assess the ability of our metric to predict the classification performance on embeddings of 3 pretrained CLIP (Radford et al., 2021) models on 4 datasets from the Breeds Hierarchy (Santurkar et al., 2020), namely entity 13, entity 30, living 17 and non living 26.
5.1 Models
For the purpose of this study we use CLIP-ViT transformers ViT-L/14, ViT-B/32, ViT-B/16 trained on 224 px $\times$ 224 px images as image encoders. Where ViT-L/14 encodes each image into a 768 dimensional embedding and ViT-B/16 and ViT-B/32 encode each image as a 512 dimensional embedding. For each of the networks we obtained model embeddings for the final image encoder as well as the mean embeddings of internal Multi-Head Self Attention Layers. We train a linear probe on the ImageNet (Deng et al., 2009) training subset of the Breeds dataset as well learn the parameters for estimating the MAP. We then correlate the performance of these 2 metrics to better understand the behaviour of our trained posterior and present the results in Figure 9.
5.2 Experimental Analysis
Table 3: In this table we show the Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(a) on the training and validation sets of entity 13 dataset of the Breeds Hierarchy.
Table 4: Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(b) on the training and validation sets of entity 30 dataset of the Breeds Hierarchy.
Table 5: Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(c) on the training and validation sets of living 17 dataset of the Breeds Hierarchy.
Entity 13
Data $\rightarrow$ Train Set Val. Set
Reg. $\downarrow$ B16 B32 L14 B16 B32 L14
$Diag$99 .99 .98 .99 .99 .97
$10^{-3}$97 .99 .97 .97 .98 .95
$10^{-6}$97 .96 .98 .99 .98 .99
$10^{-9}$96 .95 .98 .99 .99 .99
Entity 30
Data $\rightarrow$ Train Set Val. Set
Reg. $\downarrow$ B16 B32 L14 B16 B32 L14
$Diag$99 .99 .99 .99 .99 .98
$10^{-3}$98 .97 .98 .98 .98 .98
$10^{-6}$85 .83 .89 .97 .96 .99
$10^{-9}$8 .76 .87 .96 .96 .98
Living 17
Data $\rightarrow$ Train Set Val. Set
Reg. $\downarrow$ B16 B32 L14 B16 B32 L14
$Diag$99 .99 .99 .99 .99 .98
$10^{-3}$93 .93 .95 .97 .96 .99
$10^{-6}$68 .65 .72 .93 .94 .97
$10^{-9}$57 .49 .64 .93 .95 .97
Non Living 26
Data $\rightarrow$ Train Set Val. Set
Reg. $\downarrow$ B16 B32 L14 B16 B32 L14
$Diag$99 .99 .98 .99 .99 .98
$10^{-3}$98 .95 .98 .98 .98 .97
$10^{-6}$72 .67 .75 .97 .97 .99
$10^{-9}$54 .42 .64 .96 .97 .98
Table 3: In this table we show the Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(a) on the training and validation sets of entity 13 dataset of the Breeds Hierarchy.
Table 4: Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(b) on the training and validation sets of entity 30 dataset of the Breeds Hierarchy.
Table 5: Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(c) on the training and validation sets of living 17 dataset of the Breeds Hierarchy.
Table 6: Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(d) on the training and validation sets of non living 26 dataset of the Breeds Hierarchy.
For the purpose of this study we measure the correlation between MAP and linear probe accuracies learned and computed over the embeddings of the training set and validation set. The results for which are shown in Table 6, Table 6,Table 6 and Table 6 for 8(a), 8(b), 8(c) and 8(d) respectively. In Figure 9 we demonstrate the on a macro level both MAP and linear probe performance correlate. This is across various settings of regularizations for our learned metric over various Breeds Datasets and CLIP Models.
(a) CLIP Models analyzed with multiple regularizations - entity13. The pearson correlations between layerwise MAPs for various regularizations and layerwise linear probe accuracies is shown in Table 6.
(b) CLIP Models analyzed with multiple regularizations - entity30. The pearson correlations between layerwise MAPs for various regularizations and layerwise linear probe accuracies is shown in Table 6
(c) CLIP Models analyzed with multiple regularizations - living17. The pearson correlations between layerwise MAPs for various regularizations and layerwise linear probe accuracies is shown in Table 6
(d) CLIP Models analyzed with multiple regularizations - nonliving26. The pearson correlations between layerwise MAPs for various regularizations and layerwise linear probe accuracies is shown in Table 6
Figure 9: In this figure we present the evolution of layerwise MAP and linear probe accuracies for all CLIP Models on all datasets across varying regularizations strengths to demonstrate correlations between MAP and linear probe performance. The quantitative results for each settings are shown from Table 6-Table 6.
However despite this correlation, the metric is not a reliable indicator of embedding performance as many a times it can predict lower values for models which have higher linear separability as indicated by the linear probe accuracy for the embedding of that given layer, we show one such counter study in Figure 10 with additional results available in Appendix A. In 9(a) for a MAP model which assumes uncorrelated dimensions we observe that for many earlier and later layers of the CLIP models, the order of performance predicted by log posterior isn’t the same as one predict by linear probes on the embedding of that layer. This is particularly evident in case of ViT-L/14 model at both the initial and final layers of the network. A similar phenomena is also observed in 9(b), which imposes a higher regularization on the class covariance matrix learned to estimate the posterior. But this issue doesn’t completely resolve, although is less severe, when reducing the regularization on the learned class covariance as shown in 9(c) and 9(d) for regularizations $10^{-6}$ and $10^{-9}$ respectively.
(a) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 13 breeds dataset. Class covariance matrix assumes uncorrelated embedding dimensions, thereby just using the diagonal of the covariance matrix.
(b) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 13 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-3}$
(c) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 13 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-6}$
(d) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 13 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-9}$
Figure 10: Layerwise multi model analysis of log posterior performance with linear probe performance for a given regularization parameter over entity 13 Breeds Dataset.
5.3 BREEDS Datasets
Breeds Hierarchy (Santurkar et al., 2020) was by leveraging and modifying the WordNet (Miller, 1994) hierarchy for the ImageNet (Deng et al., 2009) dataset to group together semantically similar classes into 1 superclass. The original purpose of creating this hierarchy was to use the subpopulations present in superclasses to detect a model’s robustness to distribution shift but for the purpose of this study we leverage the entire datasets to evaluate the performance of our metric in predicting generalization of CLIP models on those embeddings.
6 Conclusion
This work introduces a novel method for evaluating pretrained models. Instead of using costly and time-consuming fine-tuned downstream tasks for evaluation, we propose using the consistency between entity embeddings and their associated meta features as a performance metric. Our method has been effectively tested across various domains and datasets in relational datasets, Natural Language Processing and Computer Vision, providing a more efficient and equally rigorous alternative for pretrained model evaluation.
Appendix
Appendix A Fixed regularization cross Model analysis of CLIP Embeddings on the BREEDS hierarchy
In this section we demonstrate supplementary results to the finding shown in Figure 10 regarding the dissonance between linear probe accuracies and log posterior values over the performance of some layers in the CLIP models. Figure 11 demonstrates the findings for entity 30, Figure 12 for living 17 and Figure 13 for non living 26.
(a) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 30 breeds dataset. Class covariance matrix assumes uncorrelated embedding dimensions, thereby just using the diagonal of the covariance matrix.
(b) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 30 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-3}$
(c) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 30 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-6}$
(d) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 30 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-9}$
Figure 11: Layerwise multi model analysis of log posterior performance with linear probe performance for a given regularization parameter over entity 30 Breeds Dataset.
(a) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for living 17 breeds dataset. Class covariance matrix assumes uncorrelated embedding dimensions, thereby just using the diagonal of the covariance matrix.
(b) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for living 17 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-3}$
(c) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for living 17 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-6}$
(d) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for living 17 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-9}$
Figure 12: Layerwise multi model analysis of log posterior performance with linear probe performance for a given regularization parameter over living 17 Breeds Dataset.
(a) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for non living 26 breeds dataset. Class covariance matrix assumes uncorrelated embedding dimensions, thereby just using the diagonal of the covariance matrix.
(b) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for non living 26 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-3}$
(c) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for non living 26 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-6}$
(d) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 13 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-9}$
Figure 13: Layerwise multi model analysis of log posterior performance with linear probe performance for a given regularization parameter over non living 26 Breeds Dataset.
Appendix B Analysis of CLIP Embeddings for ImageNet Subsets
In this section we continue the analysis on CLIP Image Encoder (Radford et al., 2021) embeddings on subsets of ImageNet based on superclass as defined by the Robustness library (Engstrom et al., 2019a; b; Santurkar et al., 2019; 2020). The experimental setup is explained next.
B.1 Datasets
To Test our approach we use 5 ImageNet(Deng et al., 2009) subsets using superclass labels based on WordNet(Miller, 1994) hierarchy from the robustness benchmarks (Engstrom et al., 2019a; b; Santurkar et al., 2019; 2020). The datasets are namely: big-12, geirhos-16, living-9, mixed-10, mixed-13. The dataset size statistics are included in Table 7 and the dataset label composition in Table 8.
Dataset Statistics of Robustness ImageNet Superclasses
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Training Samples 308494 40517 92632 77237 100233
Testing Samples 12005 1600 3603 3000 3903
Classes 12 16 9 10 13
Table 7: Dataset Statistics
Dataset
Class Composition
big-12
Dog (n02084071), Structure (n04341686), Bird (n01503061), Clothing (n03051540), Vehicle(n04576211), Reptile (n01661091), Carnivore (n02075296), Insect (n02159955), Instrument (n03800933), Food (n07555863), Furniture (n03405725), Primate (n02469914)
geirhos-16
Aircraft (n02686568), Bear (n02131653), Bicycle (n02834778), Bird (n01503061), Boat (n02858304), Bottle (n02876657), Car (n02958343), Cat (n02121808), Char (n03001627), Clock (n03046257), Dog (n02084071), Elephant (n02503517), Keyboard (n03614532), Knife (n03623556), Oven (n03862676), Truck (n04490091)
living-9
Dog (n02084071), Bird (n01503061), Arthropod (n01767661), Reptile (n01661091), Primate (n02469914), Fish (n02512053), Feline (n02120997), Bovid (n02401031), Amphibian (n01627424)
mixed-10
Dog (n02084071), Bird (n01503061), Insect (n02159955), Monkey (n02484322), Car (n02958343), Cat (n02120997), Truck (n04490091), Fruit (n13134947), Fungus (n12992868), Boat (n02858304)
mixed-13
Dog (n02084071), Bird (n01503061), Insect (n02159955), Furniture (n03405725), Fish (n02512053), Monkey (n02484322), Car (n02958343), Cat (n02120997), Truck (n04490091), Fruit (n13134947), Fungus (n12992868), Boat (n02858304), Computer (n03082979)
Table 8: ImageNet Superclass subsets
B.2 Experiments and Analysis
Given this setup we analyze the embeddings of the pre-trained CLIP Vision Transformers on the 2 data splits used to learn the parameters for the posterior distribution. For each network we compute the embeddings over all the datasets (train and test) shown in Table 7 and use the training set to learn a linear classifier in order to quantify the linear separability of CLIP embeddings and also learn the parameters of the posterior to assess the quality of embeddings. We plot the data and network wise posterior ($\log$) probability and the linear classifier accuracy and show the pearson correlation ($\rho$) between the 2 quantities for each respective network in Figure 14. 13(a) contain the results for the training set and the test set results are shown in 13(b).
(a) Train Posterior and linear probe plots
(b) Test Learned Posterior and linear probe plots
(c) Test Independent Posterior and linear probe plots
Figure 14: CLIP Embedding Posterior Analysis.
Training Set Metrics - ViT-B-16
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.219 -0.014 -0.064 -0.081 -0.102
Clustering Acc. Post. 95.336 99.721 98.517 98.191 97.852
Linear Probe Acc. 96.985 99.250 98.625 98.226 98.357
Table 9: ViT-B-16 Train Posterior, clustering performance and linear probe statistics
Training Set Metrics - ViT-B-32
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.271 -0.020 -0.109 -0.112 -0.137
Clustering Acc. Post. 94.209 99.595 97.471 97.477 97.105
Linear Probe Acc. 95.700 98.729 97.254 97.407 97.538
Table 10: ViT-B-32 Train Posterior, clustering performance and linear probe statistics
Training Set Metrics - ViT-L-14
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.125 -0.001 -0.012 -0.022 -0.029
Clustering Acc. Post. 97.356 99.975 99.732 99.512 99.388
Linear Probe Acc. 97.604 99.691 99.440 98.978 98.962
Table 11: ViT-L-14 Train Posterior, clustering performance and linear probe statistics
A more detailed presentation of results presented in 13(a) is shown in Table 9, Table 10 and Table 11. Further analyzing 13(a) for the training set we observe a strong correlation ($\rho$) between the $\log$ posterior probabilities and linear classifier accuracies. We also observe that for a given dataset a higher $\log$ posterior coincides with a higher linear classifier accuracy and a better clustering accuracy due to a higher posterior probability. We omitted the clustering accuracy plots for the posterior for brevity. But Table 9 - Table 17 contain those values.
Test Set Metrics - ViT-B-16
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Test Log. Post. -0.265 -0.225 -0.133 -0.168 -0.186
Clustering Acc. Post. 94.319 95.500 96.947 96.300 96.131
Linear Probe Test Acc. 97.143 99.125 98.279 98.200 98.411
Table 12: ViT-B-16 Test Posterior, clustering performance and linear probe statistics
Test Set Metrics - ViT-B-32
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.313 -0.254 -0.200 -0.207 -0.225
Clustering Acc. Post. 93.294 94.938 95.420 95.433 95.311
Linear Probe Acc. 94.494 97.688 97.197 97.833 97.771
Table 13: ViT-B-32 Test Posterior, clustering performance and linear probe statistics
Test Set Metrics - ViT-L-14
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.172 -0.180 -0.064 -0.082 -0.112
Clustering Acc. Post. 96.335 96.438 98.557 98.167 97.643
Linear Probe Acc. 97.943 99.375 99.223 99.100 99.154
Table 14: ViT-L-14 Test Posterior, clustering performance and linear probe statistics
Next, we continue the same analysis for 13(b) of the test set and its accompanying metrics in Table 12, Table 13 and Table 14. The first observation that carries over from our previous analysis on the training set is that for a given dataset, networks with higher posterior also have better accuracies, however as shown in 13(b) the correlation between those 2 sets of variables has reduced. Based on these observations we further investigate this by learning a posterior solely on the test data and comparing its generalization to the linear classifier and present the results in 13(c) and the corresponding metrics in Table 15, Table 16 and Table 17. Paying close attention to performance of ViT-B-16 and ViT-B-32 on the big-12 dataset we observe that while ViT-B-16 has more linearly separable embeddings across the training and test as shown by better linear probe accuracy in comparison to ViT-B-32. The $\log$ posterior and therefore the corresponding clustering accuracy of ViT-B-32 is higher than ViT-B-16, this presents a counter example to the earlier analysis. Further, looking at ViT-B-32 and ViT-L-14 we observe that $\log$ posterior for both networks is the same for all the datasets but the linear probe performance of those embeddings have variations. This non dependence between linear probe values and $\log$ posterior for perfectly clusterable data leads to an undefined Pearson Coefficient ($\rho$) as indicated by the value nan in 13(c).
Test (Ind.) Set Metrics - ViT-B-16
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.0004 0.000 0.000 0.000 0.000
Clustering Acc. Post. 99.992 100.000 100.000 100.000 100.000
Linear Probe Acc. 97.143 99.125 98.279 98.200 98.411
Table 15: ViT-B-16 Independently learned Test Posterior, its clustering performance and original linear probe statistics
Test (Ind.) Set Metrics - ViT-B-32
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. 0.000 0.000 0.000 0.000 0.000
Clustering Acc. Post. 100.000 100.000 100.000 100.000 100.000
Linear Probe Acc. 94.494 97.688 97.197 97.833 97.771
Table 16: ViT-B-32 Independently learned Test Posterior, its clustering performance and original linear probe statistics
Test (Ind.) Set Metrics - ViT-L-14
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. 0.000 0.000 0.000 0.000 0.000
Clustering Acc. Post. 100.000 100.000 100.000 100.000 100.000
Linear Probe Acc. 97.943 99.375 99.223 99.100 99.154
Table 17: ViT-L-14 Independently learned Test Posterior, its clustering performance and original linear probe statistics